# Gaussian fermionic embezzlement of entanglement

### v1

@@ 1 Introduction
Just like macroscopic bulk properties in the thermodynamic limit of many-body systems, some properties of entanglement only sharply emerge in systems with an unbounded number of degrees of freedom. Recent research has shown that these large-scale entanglement properties of many-body systems are closely tied to the von Neumann algebraic classification of the emergent subsystem operator algebras [1, 2, 3]. Of particular interest are those large-scale properties that only sharply manifest in the thermodynamic limit while still implying strong operational consequences for finite-size systems.
A paradigmatic example for such an effect is embezzlement of entanglement [4, 5]: Here, two agents acting on disjoint subsystems $A,B$ of the total system, can extract an arbitrary bipartite entangled state $|\Psi\rangle_{A^{\prime}B^{\prime}}\in{\mathbb{C}}^{d}\otimes{\mathbb{C}}^{d}$ on a system $A^{\prime}B^{\prime}$ from an embezzling state $|\Omega\rangle_{AB}\in{\mathcal{H}}$ to any error $\varepsilon>0$ while also perturbing the embezzling state arbitrarily little:
@@
Embezzling states require infinitely many degrees of freedom. Whether they exist is determined by the operator algebras describing subsystems $A$ and BB—in particular, by their emergent von Neumann algebraic type and subtype [2, 5]. Yet, their embezzling properties descend to finite-size versions of the same system in an approximate form [6, 7]. There are even systems where all states in the Hilbert space have the embezzling property. It was shown in [6] that ground state sectors of one-dimensional, critical, fermionic models with quadratic Hamiltonians fall into this class, which is consistent with the fact that the scaling limit of such systems should give rise to conformal field theories [8, 9, 10, 11, 12, 13, 14]. As a consequence, also systems with gapless, chiral edge modes described by non-interacting fermions, such as the $p_{x}+ip_{y}$ superconductor [15, 16], have embezzling ground states. It is plausible, but unproven, that all critical systems in one spatial dimension and all systems with a chiral edge in two spatial dimensions share this property. So far, the abstract operator-algebraic results say little about
@@
    how the achievable error $\varepsilon$ scales with the system size when considering finite-size systems.
We considerably advance the situation in this paper for ground states of quadratic, fermionic Hamiltonians, which are Gaussian (also called quasi-free) and hence fully described by their two-point correlation functions. Our main result is as follows: We precisely characterize when Gaussian fermionic states can act as (approximate) embezzling states under the restricted class of Gaussian unitary operations (those generated by quadratic fermionic Hamiltonians) and give precise error estimates. In particular, our result characterizes universal Gaussian embezzling families, which allow for embezzling arbitrary Gaussian entangled states using only Gaussian local operations. Examples comprise the ground states of the $XX$-spin chain or the transverse-field Ising model on increasing system sizes, or the above-mentioned $p_{x}+ip_{y}$ superconductors in a suitable geometry. We also show that these families allow to embezzle arbitrary entangled states if we allow for general (non-Gaussian) local unitary operations.
@@ 4 Main result
Let us discuss how to apply the result in concrete examples. Any covariance of a system with $n$ modes can have at best $1/n$-dense spectrum. Thus, optimal scaling of our bound is achieved if, for example, $K$ has eigenvalues $\lambda_{j}(K)=1-\frac{j}{n}$ with $j=1,\ldots,n$. Then $\varepsilon=1/n$ and we find $\kappa(F,G|K)\leq 11dn^{-1/4}$. Hence, as $n\rightarrow\infty$, any Gaussian state may be embezzled to arbitrary accuracy. We show in the appendix that a Gaussian state $\rho_{K}$ with $\varepsilon$-dense spectrum has entropy $H_{1}(\rho_{K})\geq\log(2)/(4\varepsilon)$, yielding extensive entropy. In the original bipartite setting, these entropies have to be interpreted as entanglement entropies. In non-interacting fermionic systems, extensive entanglement entropies of the ground states are incompatible with a local, translation-invariant Hamiltonian [17, 18, 19, 20, 21]. Therefore, this family of states does not appear naturally in the context of many-body physics.
@@
where $c_{\alpha}$ is a non-universal function of $\alpha$ [22, 23, 24]. The bound $H_{1}(\rho_{K})\geq\log(2)/(4\varepsilon)$ now implies $\varepsilon\geq O(1/\log(n))$ for a chain of length $2n$. On the other hand, the results of [6] show that the covariance has a continuous spectrum in the thermodynamic limit. Hence, the spectrum becomes arbitrarily dense as $n\to\infty$. Thus, our main result recovers the fact that ground states of non-interacting, critical fermionic systems are embezzling states, but the convergence is very slow.
4.1 From quasi-free to general embezzlement of entanglement.
We now show that an approximate Gaussian embezzling state can be used to embezzle arbitrary entangled states once we allow for general unitary operations instead of Gaussian unitary operations. The argument is simple: It has been shown in [25], that any entangled state that can embezzle arbitrary entangled states with Schmidt rank $2$ (i.e., qubit states up to local isometries) with fidelity $\mathrm{F}$ can also embezzle arbitrary pure entangled states with Schmidt rank $m$ with fidelity $\mathrm{F}_{m}^{2}\geq 1-(\log_{2}(m)^{2})(1-\mathrm{F}^{2})$. It thus suffices to show that it is possible to embezzle arbitrary pure entangled states on a pair of two-level systems, which in turn means that it is sufficient to be able to embezzle arbitrary mixed qubit states on a single system. However, up to local unitary freedom, Gaussian states of a single fermionic mode precisely correspond to arbitrary mixed qubit states, which completes the argument. [^1] If we wish to adhere to the parity superselection rule, we may encode a qubit into two fermionic modes via $\sigma_{X}\cong(a_{1}^{\dagger}-a_{1})(a_{2}^{\dagger}+a_{2})$ and $\sigma_{Z}\cong 2a_{1}^{\dagger}a_{1}-1$.
@@ 4.2 Embezzlement and number conservation
A curious property of embezzlement is that it can arbitrarily alter the local probability distribution of the total fermion number, despite the fact that passive Gaussian unitaries preserve the latter. This is possible because an embezzling system must contain many more modes than the system on which it embezzles and must have a very broad distribution for the number of fermions (otherwise it cannot have $\varepsilon$-dense spectrum for small $\varepsilon$). Thus, one may shift the mean value (or any other finite number of moments) of the distribution of the number of fermions on the embezzling system by a finite amount, while keeping the probability distribution almost unaffected as measured by total variation distance, see also [26] for a discussion of this point for other forms of embezzlement.
@@ 5 Proof of main result
was established and shown to be optimal under certain circumstances [27]. Unfortunately, (8) is insufficient for our purposes. Indeed, we prove in the appendix that
@@
which is a consequence of Wielandt’s theorem (see, e.g., [28, Theorem 9.G.1.a]), and the last inequality follows from $\lVert v\rVert_{2}\leq\sqrt{d}\lVert v\rVert_{\infty}$ for any $d$-dimensional vector $v$. We now make use of the following Lemma.
@@ 6 Conclusion
This result provides a finite-size explanation of the recent discovery that all one-dimensional critical, non-interacting fermionic systems have embezzling ground states. In the thermodynamic limit, the results of [6, 2] show that all quasi-local excitations of the ground state share the same embezzling properties. It would be desirable to show explicitly in the setting of quasi-free systems that the $\varepsilon$-dense spectrum of the covariance is stable against Gaussian unitary operations that act on a subsystem of fixed size, but may couple the modes of Alice with those of Bob. We leave this open for future work.
@@ Appendix A The general Gaussian formalism and the distance bound
Here we present, discuss, and prove our main result for the general case of non-passive (non-gauge-invariant) Gaussian fermionic states. There are various, equivalent, but slightly different formalisms to describe these states. One is Araki’s self-dual formalism, describing a system of $N$ fermionic modes using a complex $2N$-dimensional Hilbert space [29]. The other is the Majorana formalism, describing the same system using a real $2N$-dimensional Hilbert space, see for example [27]. While the former is more common in mathematical physics, the latter is more common in quantum information theory. We will therefore discuss the results in both formalisms in parallel. In contrast to the main text, we here take an operator algebraic perspective, where a quantum state is a positive, normalized functional on an operator algebra. Since we work with fermionic systems, the relevant operator algebra is the C* algebra generated by the canonical anti-commutation relations. Let $\mathfrak{h}$ be a complex single-particle Hilbert space with inner product $\left(\cdot,\cdot\right)_{\mathfrak{h}}$. We do not assume $\mathfrak{h}$ to be finite dimensional, unless it is explicitly stated. The associated algebra of canonical anti-commutation relations $\textup{CAR}(\mathfrak{h})$ is generated by the creation and annihilation operators (CAOs) $a,a^{\dagger}:\mathfrak{h}\to\textup{CAR}(\mathfrak{h})$ fulfilling the commutation relations
@@ A.3 Relation to passive Gaussian states
where 1 denotes the unit matrix. This is indeed the standard result, see for example [27].
@@ A.4 General Gaussian states and entanglement
Note that the Pfaffian is zero if $n$ is odd. Of course, one can also directly express the above relation in terms of the self-dual covariance $S$, see for example [30, Sec. 6.6].
@@ A.5 Gaussian unitary operations
Suppose that $\omega_{S}$ is pure, so that we can think of it as a vector $\Omega_{S}$ in a Fock space representation of $\textup{CAR}(\mathfrak{h})$. Then the automorphism $\alpha_{u}$ can be represented by a unitary $U$ on the Fock space if and only if $\lVert[S,u]\rVert_{2}<\infty$ [30]. This is, of course, always the case if $\mathfrak{h}$ is finite-dimensional.
@@ A.7 Distance bounds
where we used the triangle-inequality, $\lVert AB\rVert^{2}_{2}\leq\lVert A^{\dagger}A\rVert\lVert B\rVert_{2}^{2}$ and $\lVert S^{1/2}-T^{1/2}\rVert_{2}^{2}\leq\lVert S-T\rVert_{1}$ [31]. ∎
@@
To set up the proof, we briefly recall some facts from the theory of von Neumann algebras. To any quasi-free state $\omega_{S}$ on $\textup{CAR}(\mathfrak{h})$ we can associate a von Neumann algebra ${\mathcal{M}}_{S}:=\pi_{S}(\textup{CAR}(\mathfrak{h}))^{\prime\prime}$ via the GNS representation $\pi_{S}$ of $\omega_{S}$. If $\lVert\sqrt{S}-\sqrt{T}\rVert_{2}<\infty$, the states $\omega_{S},\omega_{T}$ are quasi-equivalent [29]. This implies that we can find a representation of ${\mathcal{M}}_{S}$ on a Hilbert space ${\mathcal{H}}_{S}$ (its standard representation), such that there are normalized vectors $\Omega_{S},\Omega_{T}\in{\mathcal{H}}_{S}$ in the associated positive cone of ${\mathcal{H}}_{S}$ which realize the two states and fulfill $\langle\Omega_{S}|\Omega_{T}\rangle\geq 0$ [32]. Then the transition-probability $P(\omega_{S},\omega_{T})$ is defined as $P(\omega_{S},\omega_{T}):=\langle\Omega_{S}|\Omega_{T}\rangle^{2}$. For two density operators $\rho,\sigma$ on a finite-dimensional Hilbert space, the vectors associated vectors $\Omega_{\rho},\Omega_{\sigma}$ simply correspond to canonical purifications and we have $P(\rho,\sigma)=\operatorname{Tr}(\rho^{1/2}\sigma^{1/2})^{2}$. Note that the transition probability is distinct from Uhlmann’s fidelity $\mathrm{F}(\rho,\sigma)=\lVert\rho^{1/2}\sigma^{1/2}\rVert_{1}$.
The Powers-Størmer inequality [31] implies
@@
Araki showed [29] (see also [33]) that the transition probability takes a particularly simple form for Gaussian states: [^5] In infinite dimensions, if $0\leq A\leq 1$, the determinant is defined via $\det(A):=\inf_{V}\det(A|_{V})$ where the infimum is taken over all finite-dimensional subspaces, see [31] for details.
@@ Appendix B Inadequacy of the trace-distance of covariances
Equation˜B.2 follows from Wielandt’s theorem and holds for any Schatten $p$-norm (see also [34] for a generalization to von Neumann algebras). To simplify notation, we will write $f=\lambda(F)^{\downarrow}$, $g=\lambda(G)^{\downarrow}\in{\mathbb{R}}^{d}$, $k=\lambda(K)^{\downarrow}\in{\mathbb{R}}^{m}$. By the above result, we need to show that
@@ Appendix C Entropy and density of spectrum for Gaussian states
By a slight extension of the proof of [31, Lemma 4.2], if $\delta\text{1}\leq A$ we have

### v2

@@ 1 Introduction
Just like macroscopic bulk properties in the thermodynamic limit of many-body systems, some properties of entanglement only sharply emerge in systems with an unbounded number of degrees of freedom. Recent research has shown that these large-scale entanglement properties of many-body systems are closely tied to the von Neumann algebraic classification of the emergent subsystem operator algebras [undef, undefa, undefb]. Of particular interest are those large-scale properties that only sharply manifest in the thermodynamic limit while still implying strong operational consequences for finite-size systems.
A paradigmatic example for such an effect is embezzlement of entanglement [undefc, undefd]: Here, two agents acting on disjoint subsystems $A,B$ of the total system, can extract an arbitrary bipartite entangled state $|\Psi\rangle_{A^{\prime}B^{\prime}}\in{\mathbb{C}}^{d}\otimes{\mathbb{C}}^{d}$ on a system $A^{\prime}B^{\prime}$ from an embezzling state $|\Omega\rangle_{AB}\in{\mathcal{H}}$ to any error $\varepsilon>0$ while also perturbing the embezzling state arbitrarily little:
@@
Embezzling states require infinitely many degrees of freedom. Whether they exist is determined by the operator algebras describing subsystems $A$ and BB—in particular, by their emergent von Neumann algebraic type and subtype [undefa, undefd]. Yet, their embezzling properties descend to finite-size versions of the same system in an approximate form [undefe, undeff]. There are even systems where all states in the Hilbert space have the embezzling property. It was shown in [undefe] that ground state sectors of one-dimensional, critical, fermionic models with quadratic Hamiltonians fall into this class, which is consistent with the fact that the scaling limit of such systems should give rise to conformal field theories [undefg, undefh, undefi, undefj, undefk, undefl, undefm]. As a consequence, also systems with gapless, chiral edge modes described by non-interacting fermions, such as the $p_{x}+ip_{y}$ superconductor [undefn, undefo], have embezzling ground states. It is plausible, but unproven, that all critical systems in one spatial dimension and all systems with a chiral edge in two spatial dimensions share this property. So far, the abstract operator-algebraic results say little about
@@
    how the achievable error $\varepsilon$ scales with the system size when considering finite-size systems (see, however, [undefp, undefe]),
  • (iii)
    and how generic the (approximate) embezzling property is.
We considerably advance the situation in this paper for ground states of quadratic, fermionic Hamiltonians, which are Gaussian (also called quasi-free) and hence fully described by their two-point correlation functions. Our main result is as follows: We precisely characterize when Gaussian fermionic states can act as (approximate) embezzling states under the restricted class of Gaussian unitary operations (those generated by quadratic fermionic Hamiltonians) and give precise error estimates. In particular, our result characterizes universal Gaussian embezzling families, which allow for embezzling arbitrary Gaussian entangled states using only Gaussian local operations. Our results imply that such states are generic among the pure entangled Gaussian states. Examples comprise the ground states of the $XX$-spin chain or the transverse-field Ising model on increasing system sizes, or the above-mentioned $p_{x}+ip_{y}$ superconductors in a suitable geometry. We also show that these families allow to embezzle arbitrary entangled states if we allow for general (non-Gaussian) local unitary operations.
@@ 4 Main result
Let us discuss how to apply the result in concrete examples. Any covariance of a system with $n$ modes can have at best $1/n$-dense spectrum. Thus, optimal scaling of our bound is achieved if, for example, $K$ has eigenvalues $\lambda_{j}(K)=1-\frac{j}{n}$ with $j=1,\ldots,n$. Then $\varepsilon=1/n$ and we find $\kappa(F,G|K)\leq 11dn^{-1/4}$. Hence, as $n\rightarrow\infty$, any Gaussian state may be embezzled to arbitrary accuracy. We show in the appendix that a Gaussian state $\rho_{K}$ with $\varepsilon$-dense spectrum has entropy $H_{1}(\rho_{K})\geq\log(2)/(4\varepsilon)$, yielding extensive entropy. In the original bipartite setting, these entropies have to be interpreted as entanglement entropies. In non-interacting fermionic systems, extensive entanglement entropies of the ground states are incompatible with a local, translation-invariant Hamiltonian [undefq, undefr, undefs, undeft, undefu]. Therefore, this family of states does not appear naturally in the context of many-body physics.
@@
where $c_{\alpha}$ is a non-universal function of $\alpha$ [undefv, undefw, undefx]. The bound $H_{1}(\rho_{K})\geq\log(2)/(4\varepsilon)$ now implies $\varepsilon\geq O(1/\log(n))$ for a chain of length $2n$. On the other hand, the results of [undefe] show that the covariance has a continuous spectrum in the thermodynamic limit. Hence, the spectrum becomes arbitrarily dense as $n\to\infty$. Thus, our main result recovers the fact that ground states of non-interacting, critical fermionic systems are embezzling states, but the convergence is very slow.
4.1 From quasi-free to general embezzlement of entanglement
We now show that an approximate Gaussian embezzling state can be used to embezzle arbitrary entangled states once we allow for general unitary operations instead of Gaussian unitary operations. The argument is simple: It has been shown in [undefy], that any entangled state that can embezzle arbitrary entangled states with Schmidt rank $2$ (i.e., qubit states up to local isometries) with fidelity $\mathrm{F}$ can also embezzle arbitrary pure entangled states with Schmidt rank $m$ with fidelity $\mathrm{F}_{m}^{2}\geq 1-(\log_{2}(m)^{2})(1-\mathrm{F}^{2})$. It thus suffices to show that it is possible to embezzle arbitrary pure entangled states on a pair of two-level systems, which in turn means that it is sufficient to be able to embezzle arbitrary mixed qubit states on a single system. However, up to local unitary freedom, Gaussian states of a single fermionic mode precisely correspond to arbitrary mixed qubit states, which completes the argument. [^1] If we wish to adhere to the parity superselection rule, we may encode a qubit into two fermionic modes via $\sigma_{X}\cong(a_{1}^{\dagger}-a_{1})(a_{2}^{\dagger}+a_{2})$ and $\sigma_{Z}\cong 2a_{1}^{\dagger}a_{1}-1$.
@@ 4.2 Embezzlement and number conservation
A curious property of embezzlement is that it can arbitrarily alter the local probability distribution of the total fermion number, despite the fact that passive Gaussian unitaries preserve the latter. This is possible because an embezzling system must contain many more modes than the system on which it embezzles and must have a very broad distribution for the number of fermions (otherwise it cannot have $\varepsilon$-dense spectrum for small $\varepsilon$). Thus, one may shift the mean value (or any other finite number of moments) of the distribution of the number of fermions on the embezzling system by a finite amount, while keeping the probability distribution almost unaffected as measured by total variation distance, see also [undefz] for a discussion of this point for other forms of embezzlement.
@@ 5 Proof of main result
was established and shown to be optimal under certain circumstances [undefaa]. Unfortunately, (8) is insufficient for our purposes. Indeed, we prove in the appendix that
@@
which is a consequence of Wielandt’s theorem (see, e.g., [undefab, Theorem 9.G.1.a]), and the last inequality follows from $\lVert v\rVert_{2}\leq\sqrt{d}\lVert v\rVert_{\infty}$ for any $d$-dimensional vector $v$. We now make use of the following Lemma.
@@ 6 Conclusion
This result provides a finite-size explanation of the recent discovery that all one-dimensional critical, non-interacting fermionic systems have embezzling ground states. In the thermodynamic limit, the results of [undefe, undefa] show that all quasi-local excitations of the ground state share the same embezzling properties. It would be desirable to show explicitly in the setting of quasi-free systems that the $\varepsilon$-dense spectrum of the covariance is stable against Gaussian unitary operations that act on a subsystem of fixed size, but may couple the modes of Alice with those of Bob. We leave this open for future work.
@@ Appendix A The general Gaussian formalism and the distance bound
Here we present, discuss, and prove our main result for the general case of non-passive (non-gauge-invariant) Gaussian fermionic states. There are various, equivalent, but slightly different formalisms to describe these states. One is Araki’s self-dual formalism, describing a system of $N$ fermionic modes using a complex $2N$-dimensional Hilbert space [undefac]. The other is the Majorana formalism, describing the same system using a real $2N$-dimensional Hilbert space, see for example [undefaa]. While the former is more common in mathematical physics, the latter is more common in quantum information theory. We will therefore discuss the results in both formalisms in parallel. In contrast to the main text, we here take an operator algebraic perspective, where a quantum state is a positive, normalized functional on an operator algebra. Since we work with fermionic systems, the relevant operator algebra is the C* algebra generated by the canonical anti-commutation relations. Let $\mathfrak{h}$ be a complex single-particle Hilbert space with inner product $\left(\cdot,\cdot\right)_{\mathfrak{h}}$. We do not assume $\mathfrak{h}$ to be finite dimensional, unless it is explicitly stated. The associated algebra of canonical anti-commutation relations $\textup{CAR}(\mathfrak{h})$ is generated by the creation and annihilation operators (CAOs) $a,a^{\dagger}:\mathfrak{h}\to\textup{CAR}(\mathfrak{h})$ fulfilling the commutation relations
@@ A.3 Relation to passive Gaussian states
where 1 denotes the unit matrix. This is indeed the standard result, see for example [undefaa].
@@ A.4 General Gaussian states and entanglement
Note that the Pfaffian is zero if $n$ is odd. Of course, one can also directly express the above relation in terms of the self-dual covariance $S$, see for example [undefad, Sec. 6.6].
@@ A.5 Gaussian unitary operations
Suppose that $\omega_{S}$ is pure, so that we can think of it as a vector $\Omega_{S}$ in a Fock space representation of $\textup{CAR}(\mathfrak{h})$. Then the automorphism $\alpha_{u}$ can be represented by a unitary $U$ on the Fock space if and only if $\lVert[S,u]\rVert_{2}<\infty$ [undefad]. This is, of course, always the case if $\mathfrak{h}$ is finite-dimensional.
@@ A.7 Distance bounds
where we used the triangle-inequality, $\lVert AB\rVert^{2}_{2}\leq\lVert A^{\dagger}A\rVert\lVert B\rVert_{2}^{2}$ and $\lVert S^{1/2}-T^{1/2}\rVert_{2}^{2}\leq\lVert S-T\rVert_{1}$ [undefae]. ∎
@@
To set up the proof, we briefly recall some facts from the theory of von Neumann algebras. To any quasi-free state $\omega_{S}$ on $\textup{CAR}(\mathfrak{h})$ we can associate a von Neumann algebra ${\mathcal{M}}_{S}:=\pi_{S}(\textup{CAR}(\mathfrak{h}))^{\prime\prime}$ via the GNS representation $\pi_{S}$ of $\omega_{S}$. If $\lVert\sqrt{S}-\sqrt{T}\rVert_{2}<\infty$, the states $\omega_{S},\omega_{T}$ are quasi-equivalent [undefac]. This implies that we can find a representation of ${\mathcal{M}}_{S}$ on a Hilbert space ${\mathcal{H}}_{S}$ (its standard representation), such that there are normalized vectors $\Omega_{S},\Omega_{T}\in{\mathcal{H}}_{S}$ in the associated positive cone of ${\mathcal{H}}_{S}$ which realize the two states and fulfill $\langle\Omega_{S}|\Omega_{T}\rangle\geq 0$ [undefaf]. Then the transition-probability $P(\omega_{S},\omega_{T})$ is defined as $P(\omega_{S},\omega_{T}):=\langle\Omega_{S}|\Omega_{T}\rangle^{2}$. For two density operators $\rho,\sigma$ on a finite-dimensional Hilbert space, the vectors associated vectors $\Omega_{\rho},\Omega_{\sigma}$ simply correspond to canonical purifications and we have $P(\rho,\sigma)=\operatorname{Tr}(\rho^{1/2}\sigma^{1/2})^{2}$. Note that the transition probability is distinct from Uhlmann’s fidelity $\mathrm{F}(\rho,\sigma)=\lVert\rho^{1/2}\sigma^{1/2}\rVert_{1}$.
The Powers-Størmer inequality [undefae] implies
@@
Araki showed [undefac] (see also [undefag]) that the transition probability takes a particularly simple form for Gaussian states: [^5] In infinite dimensions, if $0\leq A\leq 1$, the determinant is defined via $\det(A):=\inf_{V}\det(A|_{V})$ where the infimum is taken over all finite-dimensional subspaces, see [undefae] for details.
@@ Appendix B Inadequacy of the trace-distance of covariances
Equation˜B.2 follows from Wielandt’s theorem and holds for any Schatten $p$-norm (see also [undefah] for a generalization to von Neumann algebras). To simplify notation, we will write $f=\lambda(F)^{\downarrow}$, $g=\lambda(G)^{\downarrow}\in{\mathbb{R}}^{d}$, $k=\lambda(K)^{\downarrow}\in{\mathbb{R}}^{m}$. By the above result, we need to show that
@@ Appendix C Entropy and density of spectrum for Gaussian states
By a slight extension of the proof of [undefae, Lemma 4.2], if $\delta\text{1}\leq A$ we have